OpenStudy (keana):

Identify the range of the equation y = -x2 - 12x + 3. A:All real numbers B:y ≤ 39 C:y ≥ −6 D:y ≤ −105

OpenStudy (anonymous):

now the range huh?

OpenStudy (anonymous):

locate the vertex using x=-b/(2a)... what value did you get?

OpenStudy (keana):

if i remembered how to do that, this wouldn't be an issue

OpenStudy (anonymous):

use x=-b/(2a) can you identify a, b in \(\large y=ax^2+bx+c \) according to your equation?

OpenStudy (keana):

i really don't have a clue how to do it, which is why i need it to be explained !

OpenStudy (anonymous):

here all you have to do is match up with the equation i gave u... \(\large y=ax^2+bx+c \) \(\large y=-x^2-12x+3 \) just match up the numbers so you can find out what a and b are...

OpenStudy (keana):

ok but i don't think i understand what that does. how does it change anything?? isn't it still the same??? i need an explanation please because i don't understand

OpenStudy (anonymous):